# The ambiguous AT2022rze: Changing-look AGN mimicking a supernova in a merging galaxy system

### v1

@@ 3 Analysis
AT2022rze was originally classified as a SN IIn. Because of its peak absolute magnitude, it is included in the SLSN II sample presented by Pessi et al. (2024). AT2022rze does not stand out from the rest of the SLSN II sample. However, Pessi et al. (2024) only include observations up to December 12th 2022, an arbitrary date that could result in missed light curve characteristics. This event was re-classified as an AGN in December 16th 2024. Thus, we go back to AT2022rze and find that its light curve continues to show activity for over 700 days after peak. Yet, there is an apparent lack of evolution of the spectral lines that resulted in the original SN classification, suggesting that such classification could still be valid. In the following, we analyze the available observations in order to decipher the true nature of this transient.
@@ 3.1.1 Light curve peak and timescales
The gaps in observations around the first peak prevent an accurate estimate of the peak epoch in any given filter. Given the overlap between the and filters, we consider no color evolution around the time of peak, and include the observations in all these bands to estimate an overall approximate peak epoch. To do this, we use Gaussian Process (GP, e.g: Rasmussen & Williams, 2006) to interpolate the observations that cover the main (more luminous) light curve peak in the mentioned filters. GP is implemented utilizing the GPy Python package101010https://gpy.readthedocs.io/en/deploy/.. We consider the epoch and absolute magnitude of the peak of the GP median to be the peak epoch and peak absolute magnitude of AT2022rze. Errorbars are estimated from the associated confidence interval. The interpolation together with the confidence region is shown as dashed pink lines in the bottom panel of Fig. 2. The adopted peak epoch is MJD 59836.5, and the adopted peak absolute magnitude is 20.2 0.2 mag, shown as a pink star in the bottom panel of Fig. 2. Since the interpolation considers and , this peak epoch does not correspond to a single band, but we would associate the peak with the band, as this is the filter with the wider passband and the most observations in the considered region. All the phases in this work are presented as rest frame days with respect to the adopted peak, unless stated otherwise.
@@
Pessi et al. (2024) follows the work of Chen et al. (2023) and define rise and decline timescales as flux fractions with respect to peak. The considered fractions are 10% and 1/e, which correspond to mag 2.5 and mag 1.09, respectively. Dashed gray horizontal lines in Fig. 2 indicate the considered mag’s. Using the GP interpolation described above, we can estimate a 1/e rise time t days. trise,10% cannot be determined due to the lack of early observations. We can see that the light curve shows many wiggles and bumps after peak, it also has several gaps in the observations. Thus, GP interpolation is not ideal to interpolate the whole light curve. Instead, we interpolate the post-peak band light curve using the Automated Loess Regression (ALR) pipeline presented by Rodríguez et al. (2019). The resulting interpolation is presented as a gray-dashed line in Fig. 2. We use this ALR interpolation to determine the decline timescales tdec,1/e and tdec,10%, which are and days, respectively.
@@
These rise times of AT2022rze are comparable to the rise times presented for SLSNe II by Pessi et al. (2024). They report a median trise,1/e of 33.8, 37.0 and 46.9 days in the bands, respectively. On the contrary, the decline times of AT2022rze are much longer than those presented for SLSNe II. The median tdec,1/e of SLSNe II being 75.0, 89.3 and 98.3 days in the bands, respectively; and the median tdec,10% being 235.8, 248.1, 246.8 days in the bands, respectively. However, Pessi et al. (2024) find that a fraction of slow declining SLSNe II exist ( 5% of their sample), that take over a year to fade.
@@ 3.1.2 Color evolution
The full ZTF band light curves of AT2022rze were interpolated with ALR. These interpolations were used to construct the color curve presented in Fig. 4. We compare the observed color of AT2022rze to those of the SLSN II sample presented by Pessi et al. (2024) and to those of TDEs presented by Yao et al. (2023). We find better agreement with SLSNe II (although see Sect. 3.3). We note that we do not have color information before or around peak, but the evolution at the earliest observed phases seem to follow the trend of the SLSN II sample.
@@ 3.2 Spectra
The spectral evolution of AT2022rze is presented in Fig. 3. The first available spectrum has low resolution and was observed 58 days after peak. The following observed spectra seem to show little to no spectral line evolution. In Fig. 6 we show the identified spectral lines on the spectrum obtained at 173 d. Besides the hydrogen lines that originally led to the Type II SN classification, we can see several lines from He i. In particular, the He i 5876 is blended with Na id. It is unclear whether Fe ii lines are present at 4924 and 5018, as they overlap with He i 4922 and 5016 respectively. There is also weak S ii 6715 and possibly C ii 9234, although the redder part of the spectrum is quite noisy, making it difficult to properly identify lines. We see O i 8446 but there is barely any indication of O i 7774. The absence of O i 7774 and presence of O i 8446 could indicate the presence of Bowen resonance-fluorescence mechanisms, which may indicate the existence of intense far-UV radiation fields (see Osterbrock & Ferland, 2006). We tentatively identify He ii 4686. In addition, we see several high excitation lines, such as [Ne v] 3426, [Ne iii] 3869, [Fe vii] 3759, [Fe v] 4180.6, [Ni xii] 4232, [O iii] 4363, [Ar xiv] 4412, [O iii] 4959, [O iii] 5007, [Fe vi] 5176, [Fe vii] 5276, [Fe xiv] 5303, [Ar x] 5533, [Fe vii] 5721, [Fe vii] 6087, [Fe x] 6374, and [Fe xi] 7892. Some of these lines are typically known as coronal lines, and require intense UV/X-ray radiation or collisionally ionized hot gas to be formed (see Osterbrock & Ferland, 2006). The spectra show little evolution and the same lines can be identified during most of the spectral sequence.
@@ 3.3 Host
AT2022rze has been ingested by the Blast web application111111https://blast.scimma.org/ (Jones et al., 2024), which provides global and local (within 2 kpc) properties of a transients’ host galaxy. The software uses UV, optical and IR archival data of the host to fit the spectral energy distribution (SED) and estimate different characteristics. Blast considers the (global) host galaxy of AT2022rze to be enclosed by the green ellipse shown in the top right panel of Fig. 1. Throughout this work, we have been considering that to be the host of AT2022rze, WISEAJ122252.07760253.9, which we assume to be a single galaxy.
Images from many shallow surveys (including WISE) suggest that host WISEAJ122252.07760253.9 has an irregular morphology. A visual inspection of deeper images from the Legacy Survey suggest that this is because it is formed by two separate galaxies undergoing a merging event (see Fig. 1). To explore this possibility and better understand the transient’s location within its host, we perform a qualitative analysis of the Legacy Survey band image shown in the top right panel of Fig. 1. We first displayed the image on the SAOImage DS9 software (Smithsonian Astrophysical Observatory, 2000; Joye & Mandel, 2003), considering an intensity range from 0.1 to 0.6. Then, we used the contours analysis tool to find the curves defined by constant value pixels, and find two different regions of maximum intensity. Thus, we conclude that indeed, the host is a merging system.
@@
The location of AT2022rze is then offset 4.8(7.1 kpc), 5.4(8.1 kpc) and 0.1(0.2 kpc) from the center of the host, the center of galaxy A and the center of galaxy B, respectively. This is a rough approximation and dedicated analysis of deeper, higher resolution images is needed to better locate the transient. Still, the host galaxy appears to be a galaxy merger with disturbed morphology and AT2022rze appears coincident with one of the nuclei of this galaxy merger (identified as galaxy B).
The Blast analysis concludes that the host of AT2022rze is consistent with an AGN, both in their global and local properties, even though the latter only include optical bands. Note that the global analysis includes the entirety of the host while the local analysis only includes galaxy B (see upper right panel of Fig. 1). While the Blast approach has some limitations (see Jones et al., 2024), it offers a useful and efficient means of rapidly assessing the host galaxy and estimating its properties. The parameters inferred for AT2022rze can be found at https://blast.scimma.org/transients/2022rze/. Most relevant to decipher the nature of AT2022rze are: the estimated median mass, which is 5.5 M⊙ for the global analysis, and 4.9 M⊙ for the local analysis; the estimated star formation rate (SFR), which is 4.3 M⊙ y-1 and 0.03 M⊙ y-1 for the global and local analysis, respectively; and the AGN role (or , fraction of bolometric luminosity emission due to AGN), which is 1.9% and 18% for the global and local analysis, respectively. These parameters indicate that the host is a massive galaxy with a moderate star formation and a very weak influence of an AGN. On the contrary, the local analysis is consistent with a low-mass galaxy (similar to a dwarf galaxy), with very low star formation and a non-negligible probability of AGN activity. Although this analysis is limited and insufficient to securely claim the presence of an AGN, the differences between the global and local analysis, together with the qualitative assessment of the morphology of the galaxy, indicated that galaxy B, at which AT2022rze is centered, is indeed a smaller galaxy with a nuclear AGN merging with a bigger, star forming galaxy.
@@
where is the sum of the pEW of both Na id components (Poznanski et al., 2012). This method has several caveats and becomes insensitive at pEW(Na id) 1Å (Phillips et al., 2013). From the spectrum of AT2022rze, we obtain Å and Å this is consistent with the pEW of the same lines in the galaxy spectra. Thus, attempts to estimate host extinction using this method would be inadequate. Phillips et al. (2013) also consider the diffuse interstellar band (DIB) 5780 and the K i lines 7665 and 7699 as proxy for independently estimating host extinction. We do not clearly see these lines neither on the transient nor on the galaxies (see Fig. 8). Still, we measure a Balmer decrements at all phases (see Fig. 12 and Table 7), which is typically indicative of extinction.
Given that there is evidence of host extinction, we decided to consider equation 2, even if inadequate, to provide a rough estimate of the possible impact of such extinction. We obtained a host extinction of E mag that, considering R, implies an A mag. This would imply a peak absolute magnitude of 24.5 mag ( 4 mag brighter than without considering host extinction) and total radiated energies of erg (considering the pseudo bolometric peak, see Sect. 3.1.3), similar to the radiated energy of other ambiguous transients (e.g., Kankare et al., 2017; Mattila et al., 2018).
The color of AT2022rze also significantly changes when considering host extinction. The observed color is presented in the left panel of Fig. 4, while the absolute magnitude and light curves are presented in the right panel of the same figure. Inspecting the light curves before applying host extinction (circles), we would conclude that AT2022rze is rather red, and definitely redder than TDEs at early times. However, after applying the estimated host extinction, the transient (stars) seems much bluer. Given our approximations, the true host galaxy extinction would likely be somewhere between these two extremes. In order to effectively determine the host extinction, we would need better resolution spectra, to be able to identify and measure the column density of the extinction indicators. Still, our rough estimation suggest that AT2022rze may be bluer than observed.
@@ 3.4 Comparison to other events
Although AT2022rze was classified as a SN and shows some SN characteristics, it was later classified as an AGN. Yet, both classifications are based on spectral characteristics, which highlights the uncertainties present when performing spectral matching. When the classification between SN and AGN is dubious, one should also consider the possibility of a TDE, as these nuclear transients can also appear similar. Below we compare AT2022rze to well studied events, including: SNe (SN 2010jl, Fransson et al. 2014; iPTF13z, Nyholm et al. 2017; iPTF14hls, Arcavi et al. 2017; Sollerman et al. 2019), AGNs (Mrk1044 and ESO424-12 both obtained from the BAT AGN Spectroscopic Survey121212BASS: https://www.bass-survey.com/) and TDEs (SDSS1115+0544, originally considered a CLAGN by Yan et al. 2019 it was re-classified as a TDE by Wang et al. 2022 and Zhang et al. 2025; ASASSN-14li, Holoien et al. 2016; Charalampopoulos et al. 2022) that show similarities, either in the light curve or spectra, to AT2022rze.
@@
These three lines are of particular interest because, together with [N ii] 6584, can be used as AGN diagnostic when plotting corresponding ratios in a BPT diagram (Baldwin et al., 1981). We caution that [N ii] 6584 is not clearly visible in the spectra of AT2022rze, even though it is present in the spectra of the host (see Fig. 7). We assume that the [N ii] lines are blended with H and use the the Python package lmfit (Newville et al., 2024) to fit the H using a three Gaussian model. We center one of the Gaussians at the H rest wavelength and the other two at the rest wavelengths of [N ii] 6548 and 6584, respectively. We consider the sigma of [N ii] 6584 to be the same of the [O iii] 5007 line (as forbidden lines should have similar dispersion). The results are presented in Fig. 11. While BPT diagrams are known to suffer from biases related to H distribution within galaxies (e.g., Trump et al., 2015) and their applicability to transient phenomena remains uncertain (e.g., Frederick et al., 2019), they can nonetheless provide useful additional context when interpreting events. The measurements for AT2022rze span a broad range that includes the regions corresponding both to star forming galaxies and liners, the latter formed both by weak AGNs and galaxies that no longer form stars. All the relevant lines are clearly visible in the host spectra, and measurements are straightforward. Galaxy A (grey markers) is consistent with a star forming galaxy, while galaxy B is consistent with an AGN.
@@
WISE colors can also be used as diagnostic to determine whether an event may be an (obscure) AGN (see Wright et al., 2010b). In particular, a W1W2 0.5 mag could suggest an AGN origin if it is accompanied by a W2W3 2.5 mag. The AllWISE W1W2 color measured near the position of AT2022rze before it went off are < 0.5 mag (see Table 1). In particular, the color corresponding to the position closer to the transient (labeled “other” in Table 1) is W1W2 mag. This value is similar to the colors measured for CLAGNs (Frederick et al., 2021). That being said, the WISE light curve luminosity (see Fig. 2) is comparable to the luminosity observed in other CLAGNs and ANTs (López-Navas et al., 2023; Hinkle, 2024). ANTs are ambiguous transients associated to either CLAGNs or obscured TDEs (Wiseman et al., 2025). Distinguishing between the two is challenging, and some previously classified CLAGNs have been reclassified as TDEs, as is the example of SDSS1115+0544 (Yan et al., 2019; Wang et al., 2022; Zhang et al., 2025).
Grotova et al. (2025) present a sample of TDEs detected with eROSITA. The eROSITA dectection of AT2022rze is fainter than their TDE sample. However, their W1W2 cut to remove AGNs is similar to that measured for AT2022rze. Charalampopoulos et al. (2022) perform a detailed spectral analysis of TDEs. In Fig. 12 we compare the H to H ratio and the H FWHM velocity of AT2022rze to their sample. The corresponding measured parameters are listed in Table 7. In the top panel of Fig. 12 we can see that the H to H ratio of AT2022rze are among the highest values measured for TDEs. AT2022rze also shows a progressive increase in this parameter over time that is not so obvious for TDEs. We measured the H FWHM velocity of AT2022rze by fitting Gaussian models with lmfit. The associated error bars were calculated by arbitrarily modifying the limits of the line continua and repeating the Gaussian fit. In the bottom panel of Fig. 12 we can see that the measured velocity for AT2022rze is much lower than the velocities measured for the considered sample of TDEs.
@@
TDEs are associated to massive black holes (MBHs) with masses M⊙ (although higher masses are possible for highly-spinning black holes and low-density stars Kesden, 2012; Leloudas et al., 2016). Yao et al. (2023) present a relation between the host galaxy stellar mass and the mass of the MBH producing the TDE. Considering the local galaxy stellar mass presented in Sect. 3.3, we infer a BH mass M⊙. Considering the relation between BH mass and galaxy stellar mass presented by Reines & Volonteri (2015), we find M⊙. These values indicate that a TDE is possible. While not conclusive, these values are much lower than those presented by Reines & Volonteri (2015) for their AGN sample. However, the population of obscured AGN associated to lower-mass BHs is increasing (e.g Purohit et al., 2025)

### v2

@@ 3 Analysis
AT2022rze was originally classified as a SN IIn. Because of its peak absolute magnitude, it is included in the SLSN II sample presented by 2025A&A...695A.142P. AT2022rze does not stand out from the rest of the SLSN II sample. However, 2025A&A...695A.142P only include observations up to December 12th 2022, an arbitrary date that could result in missed light curve characteristics. This event was re-classified as an AGN in December 16th 2024. Thus, we go back to AT2022rze and find that its light curve continues to show activity for over 700 days after peak. Yet, there is an apparent lack of evolution of the spectral lines that resulted in the original SN classification, suggesting that such classification could still be valid. In the following, we analyze the available observations in order to decipher the true nature of this transient.
@@ 3.1.1 Light curve peak and timescales
The gaps in observations around the first peak prevent an accurate estimate of the peak epoch in any given filter. Given the overlap between the and filters, we consider no color evolution around the time of peak, and include the observations in all these bands to estimate an overall approximate peak epoch. To do this, we use Gaussian Process (GP, e.g: 2006gpml.book.....R) to interpolate the observations that cover the main (more luminous) light curve peak in the mentioned filters. GP is implemented utilizing the GPy Python package101010https://gpy.readthedocs.io/en/deploy/.. We consider the epoch and absolute magnitude of the peak of the GP median to be the peak epoch and peak absolute magnitude of AT2022rze. Errorbars represent the standard deviation of the estimated peak of the median of 1000 posterior samples. The interpolation together with the confidence region is shown as dashed pink lines in the bottom panel of Fig. 2. The adopted peak epoch is MJD 59836.5 2.4, and the adopted peak absolute magnitude is 20.2 0.2 mag, shown as a pink star in the bottom panel of Fig. 2. Since the interpolation considers and , this peak epoch does not correspond to a single band, but we would associate the peak with the band, as this is the filter with the wider passband and the most observations in the considered region. All the phases in this work are presented as rest frame days with respect to the adopted peak, unless stated otherwise.
@@
2025A&A...695A.142P follows the work of 2023ApJ...943...41C and define rise and decline timescales as flux fractions with respect to peak. The considered fractions are 10% and 1/e, which correspond to mag 2.5 and mag 1.09, respectively. Dashed gray horizontal lines in Fig. 2 indicate the considered mag’s. Using the GP interpolation described above, we can estimate a 1/e rise time t days. trise,10% cannot be determined due to the lack of early observations. We can see that the light curve shows many wiggles and bumps after peak, it also has several gaps in the observations. Thus, GP interpolation is not ideal to interpolate the whole light curve. Instead, we interpolate the post-peak band light curve using the Automated Loess Regression (ALR) pipeline presented by 2019MNRAS.483.5459R. The resulting interpolation is presented as a gray-dashed line in Fig. 2. We use this ALR interpolation to determine the decline timescales tdec,1/e and tdec,10%, which are and days, respectively.
@@
These rise times of AT2022rze are comparable to the rise times presented for SLSNe II by 2025A&A...695A.142P. They report a median trise,1/e of 33.8, 37.0 and 46.9 days in the bands, respectively. On the contrary, the decline times of AT2022rze are much longer than those presented for SLSNe II. The median tdec,1/e of SLSNe II being 75.0, 89.3 and 98.3 days in the bands, respectively; and the median tdec,10% being 235.8, 248.1, 246.8 days in the bands, respectively. However, 2025A&A...695A.142P find that a fraction of slow declining SLSNe II exist ( 5% of their sample), that take over a year to fade.
@@ 3.1.2 Color evolution
The full ZTF band light curves of AT2022rze were interpolated with ALR. These interpolations were used to construct the color curve presented in Fig. 4. We compare the observed color of AT2022rze to those of the SLSN II sample presented by 2025A&A...695A.142P and to those of TDEs presented by 2023ApJ...955L...6Y. We find better agreement with SLSNe II (although see Sect. 3.3). We note that we do not have color information before or around peak, but the evolution at the earliest observed phases seem to follow the trend of the SLSN II sample.
@@ 3.2 Spectra
The spectral evolution of AT2022rze is presented in Fig. 3. The first available spectrum has low resolution and was observed 58 days after peak. The following observed spectra seem to show little to no spectral line evolution. In Fig. 6 we show the identified spectral lines on the spectrum obtained at 173 d. Besides the hydrogen lines that originally led to the Type II SN classification, we can see several lines from He i. In particular, the He i 5876 is blended with Na id. It is unclear whether Fe ii lines are present at 4924 and 5018, as they overlap with He i 4922 and 5016 respectively. There is also weak S ii 6715 and possibly C ii 9234, although the redder part of the spectrum is quite noisy, making it difficult to properly identify lines. We see O i 8446 but there is barely any indication of O i 7774. The absence of O i 7774 and presence of O i 8446 could indicate the presence of Bowen resonance-fluorescence mechanisms, which may indicate the existence of intense far-UV radiation fields (see 2006agna.book.....O). We tentatively identify He ii 4686. In addition, we see several high excitation lines, such as [Ne v] 3426, [Ne iii] 3869, [Fe vii] 3759, [Fe v] 4180.6, [Ni xii] 4232, [O iii] 4363, [Ar xiv] 4412, [O iii] 4959, [O iii] 5007, [Fe vi] 5176, [Fe vii] 5276, [Fe xiv] 5303, [Ar x] 5533, [Fe vii] 5721, [Fe vii] 6087, [Fe x] 6374, and [Fe xi] 7892. Some of these lines are typically known as coronal lines, and require intense UV/X-ray radiation or collisionally ionized hot gas to be formed (see 2006agna.book.....O). The spectra show little evolution and the same lines can be identified during most of the spectral sequence.
@@ 3.3 Host
AT2022rze has been ingested by the Blast web application111111https://blast.scimma.org/ (2024arXiv241017322J), which provides global and local (within 2 kpc) properties of a transients’ host galaxy. The software uses UV, optical and IR archival data of the host to fit the spectral energy distribution (SED) and estimate different characteristics. Blast considers the (global) host galaxy of AT2022rze to be enclosed by the green ellipse shown in the top right panel of Fig. 1. Throughout this work, we have been considering that to be the host of AT2022rze, WISEAJ122252.07760253.9.
Images from many shallow surveys (including WISE) suggest that host WISEAJ122252.07760253.9 has an irregular morphology. A visual inspection of deeper images from the Legacy Survey suggest that this is because it is formed by two separate galaxies undergoing a merging event (see Fig. 1). To explore this possibility and better understand the transient’s location within its host, we perform a qualitative analysis of the Legacy Survey band image shown in the top right panel of Fig. 1. We first displayed the image on the SAOImage DS9 software (2000ascl.soft03002S; 2003ASPC..295..489J), considering an intensity range from 0.1 to 0.6. Then, we used the contours analysis tool to find the curves defined by constant value pixels, and find two different regions of maximum intensity, indicating that the host is a merging system.
@@
The location of AT2022rze is then offset 4.8(7.1 kpc), 5.4(8.1 kpc) and 0.1(0.2 kpc) from the center of the host, the center of galaxy A and the center of galaxy B, respectively. This is a rough approximation and dedicated analysis of deeper, higher resolution images are needed to better locate the transient. Still, the host galaxy appears to be a galaxy merger with disturbed morphology and AT2022rze appears coincident with one of the nuclei of this galaxy merger (identified as galaxy B).
The Blast analysis concludes that the host of AT2022rze is consistent with an AGN, both in their global and local properties, even though the latter only include optical bands. Note that the global analysis includes the entirety of the host while the local analysis only includes galaxy B (see upper right panel of Fig. 1). While the Blast approach has some limitations (see 2024arXiv241017322J), it offers a useful and efficient means of rapidly assessing the host galaxy and estimating its properties. The parameters inferred for AT2022rze can be found at https://blast.scimma.org/transients/2022rze/. Most relevant to decipher the nature of AT2022rze are: the estimated median mass, which is 5.5 M⊙ for the global analysis, and 4.9 M⊙ for the local analysis; the estimated star formation rate (SFR), which is 4.3 M⊙ y-1 and 0.03 M⊙ y-1 for the global and local analysis, respectively; and the AGN role (or , fraction of bolometric luminosity emission due to AGN), which is 1.9% and 18% for the global and local analysis, respectively. These parameters indicate that the host is a massive galaxy with a moderate star formation and a very weak influence of an AGN. On the contrary, the local analysis is consistent with a low-mass galaxy (similar to a dwarf galaxy), with very low star formation and a non-negligible probability of AGN activity. Although this analysis is limited and insufficient to securely claim the presence of an AGN, and the local photometry would need to be deconvolved to accurately estimate the mass of both galaxies, the differences between the global and local analysis, together with the qualitative assessment of the morphology of the galaxy, indicates that galaxy B, at which AT2022rze is centered, is indeed a smaller galaxy with a nuclear AGN merging with a bigger, star forming galaxy.
@@
where is the sum of the pEW of both Na id components (2012MNRAS.426.1465P). This method has several caveats and becomes insensitive at pEW(Na id) 1Å (2013ApJ...779...38P). From the spectrum of AT2022rze, we obtain Å and Å this is consistent with the pEW of the same lines in the galaxy spectra. Thus, attempts to estimate host extinction using this method would be inadequate. 2013ApJ...779...38P also consider the diffuse interstellar band (DIB) 5780 and the K i lines 7665 and 7699 as proxy for independently estimating host extinction. We do not clearly see these lines neither on the transient nor on the galaxies (see Fig. 8). Still, we measure a Balmer decrements at all phases (see Fig. 12 and Table 7), which is typically indicative of extinction.
Given that there is evidence of host extinction, we decided to consider equation 2, even if inadequate, to provide a rough estimate of the possible impact of such extinction. We obtained a host extinction of E mag that, considering R, implies an A mag. This would imply a peak absolute magnitude of 24.5 mag ( 4 mag brighter than without considering host extinction) and total radiated energies of erg (considering the pseudo bolometric peak, see Sect. 3.1.3), similar to the radiated energy of other ambiguous transients (e.g., 2017NatAs...1..865K; 2018Sci...361..482M).
The color of AT2022rze also significantly changes when considering host extinction. The observed color is presented in the left panel of Fig. 4, while the absolute magnitude and light curves are presented in the right panel of the same figure. Inspecting the light curves before applying host extinction (circles), we would conclude that AT2022rze is rather red, and definitely redder than TDEs at early times. However, after applying the estimated host extinction, the transient (stars) becomes much bluer. Given our approximations, the true host galaxy extinction would likely be somewhere between these two extremes. In order to effectively determine the host extinction, we would need better resolution spectra, to be able to identify and measure the column density of the extinction indicators. Still, our rough estimation suggests that AT2022rze may be bluer than observed.
@@ 3.4 Comparison to other events
Although AT2022rze was classified as a SN and shows some SN characteristics, it was later classified as an AGN. Yet, both classifications are based on spectral characteristics, which highlights the uncertainties present when performing spectral matching. When the classification between SN and AGN is dubious, one should also consider the possibility of a TDE, as these nuclear transients can also appear similar. Below we compare AT2022rze to well studied events, including: SNe (SN 2010jl, 2014ApJ...797..118F; iPTF13z, 2017A&A...605A...6N; iPTF14hls, 2017Natur.551..210A; 2019A&A...621A..30S), AGNs (Mrk1044 and ESO424-12 both obtained from the BAT AGN Spectroscopic Survey121212BASS: https://www.bass-survey.com/) and TDEs (SDSS1115+0544, originally considered a CLAGN by 2019ApJ...874...44Y it was re-classified as a TDE by 2022ApJS..258...21W and 2025arXiv250108812Z; ASASSN-14li, 2016MNRAS.455.2918H; 2022A&A...659A..34C) that show similarities, either in the light curve or spectra, to AT2022rze. We also include comparisons to samples of (SL)SN IIn (2013A&A...555A..10T; 2025A&A...695A..29S), TDEs (2022A&A...659A..34C) and CLAGNs (2024ApJ...966..128W; 2024ApJ...966...85Z).
@@
These three lines are of particular interest because, together with [N ii] 6584, can be used as AGN diagnostic when plotting corresponding ratios in a BPT diagram (1981PASP...93....5B). We caution that [N ii] 6584 is not clearly visible in the spectra of AT2022rze, even though it is present in the spectra of the host (see Fig. 7). We assume that the [N ii] lines are blended with H and use the the Python package lmfit (2021zndo....598352N) to fit the H using a three Gaussian model. We center one of the Gaussians at the H rest wavelength and the other two at the rest wavelengths of [N ii] 6548 and 6584, respectively. We consider the sigma of [N ii] 6584 to be the same of the [O iii] 5007 line (as forbidden lines should have similar dispersion). The results are presented in Fig. 11. While BPT diagrams are known to suffer from biases related to H distribution within galaxies (e.g., 2015ApJ...811...26T) and their applicability to transient phenomena remains uncertain (e.g., 2019ApJ...883...31F), they can nonetheless provide useful additional context when interpreting events. The measurements for AT2022rze span a broad range that includes the regions corresponding both to star forming galaxies and liners, the latter formed both by weak AGNs and galaxies that no longer form stars. All the relevant lines are clearly visible in the host spectra, and measurements are straightforward. Galaxy A (grey markers) is consistent with a star forming galaxy, while galaxy B is consistent with an AGN.
@@
WISE colors can also be used as diagnostic to determine whether an event may be an (obscured) AGN (see 2010AJ....140.1868W). In particular, a single epoch W1W2 0.5 mag color could suggest an AGN origin if it is accompanied by a W2W3 2.5 mag. The AllWISE W1W2 color measured near the position of AT2022rze before it went off are < 0.5 mag (see Table 1). In particular, the color corresponding to the position closer to the transient (labeled “other” in Table 1) is W1W2 mag. This value is similar to the colors measured for CLAGNs (2021ApJ...920...56F). That being said, the WISE light curve luminosity (see Fig. 2 and Sect. 4.3 for further discussion) is comparable to the luminosity observed in other CLAGNs and ANTs (2023MNRAS.524..188L; 2024MNRAS.531.2603H). ANTs are ambiguous transients associated to either CLAGNs or obscured TDEs (2025MNRAS.537.2024W). Distinguishing between the two is challenging, and some previously classified CLAGNs have been reclassified as TDEs, as is the example of SDSS1115+0544 (2019ApJ...874...44Y; 2022ApJS..258...21W; 2025arXiv250108812Z).
@@
2025A&A...697A.159G present a sample of TDEs detected with eROSITA. The eROSITA dectection of AT2022rze is fainter than their TDE sample. However, their W1W2 cut to remove AGNs is similar to that measured for AT2022rze. TDEs are associated to massive black holes (MBHs) with masses M⊙ (although higher masses are possible for highly-spinning black holes and low-density stars 2012PhRvD..85b4037K; 2016NatAs...1E...2L). 2023ApJ...955L...6Y present a relation between the host galaxy stellar mass and the mass of the MBH producing the TDE. Considering the local galaxy stellar mass presented in Sect. 3.3, we infer a BH mass M⊙. Considering the relation between BH mass and galaxy stellar mass presented by 2015ApJ...813...82R, we find M⊙. These values indicate that a TDE is possible. While not conclusive, these values are much lower than those presented by 2015ApJ...813...82R for their AGN sample. However, the population of obscured AGN associated to lower-mass BHs is increasing (e.g 2025MNRAS.538.2116P).
Fig. 12 shows a comparison of the H to H ratio (top panel) and H FWHM velocity (bottom panel) of AT2022rze to samples of TDEs, (SL)SNe IIn and CLAGNs. The H FWHM velocity of AT2022rze was obtained by fitting Gaussian models with lmfit. The associated error bars were calculated by arbitrarily modifying the limits of the line continua and repeating the Gaussian fit. The corresponding measured parameters for AT2022rze are listed in Table 7. In this case, we only focus on comparison events studied in bulk, as they are analyzed using consistent methods and are likely to involve fewer biases than a collection of results obtained from individual object analyses. We selected the TDE sample of 2022A&A...659A..34C, the sample of SN IIn of 2013A&A...555A..10T, the sample of strongly interacting (SI-) SN IIn of 2025A&A...695A..29S that includes two SLSN IIn131313We consider SLSN II any Type II event whose light curve reach a peak magnitude brighter than mag (2025A&A...695A.142P)., and the CLAGN samples of 2024ApJ...966..128W for H to H ratio and 2024ApJ...966...85Z for H FWHM velocity. The relevant parameters of CLAGNs are taken from two separate studies, as we did not find a single work that provides both. Furthermore, we show only the highest and lowest measurements of each parameter as a shaded region, since the considered studies do not account for different light-curve phases. Note that the considered phases also vary from study to study, we present rest frame days since peak similar to 2022A&A...659A..34C, although they consider peaks at different bands. 2013A&A...555A..10T consider days since discovery and 2025A&A...695A..29S days since explosion. Thus a phase shift of up to 90 days (largest rise time upper limit in 2025A&A...695A..29S) may need to be consider for the SN IIn comparison samples.
In the top panel of Fig. 12 we can see that the H to H ratios of AT2022rze are among the highest values measured for TDEs and SN IIn, and are consistent with the values measured for SI-SN IIn at later times. The comparison to CLAGN is somewhat unfair as our measurements consider the entire observed spectrum and the CLAGN measurements of 2024ApJ...966..128W subtract the galaxy component, so they represent only a lower limit in this case. In the bottom panel of Fig. 12 we can see that the measured H velocity for AT2022rze is much lower than the velocities measured for the considered sample of TDEs. At early times, the velocity of AT2022rze is consistent with the velocity of the narrow components of SN IIn and SI-SN IIn. Both 2013A&A...555A..10T and 2025A&A...695A..29S present measurements for the broad and narrow components of the H features, if these are distinguishable. We only consider the velocity of the narrow component of the sample of 2013A&A...555A..10T as it provides a good enough fit to the features they present. The sample of 2025A&A...695A..29S show significantly broader features so we include both the narrow and broad components separately. At late times the broad H component velocity of SI-SN IIn is comparable to that of AT2022rze, while the rest of the events show slower narrow component velocities, except for SN 2021acya. The CLAGNs’ lower H velocities are also consistent with the late time velocities of AT2022rze.